OpenStudy (anonymous):

a company builds computers. it cost $6,700 to build 10 computers and 12,200 to build 20 computers. Which equation models the cost, C(x), as a linear function of the number of computers built,x?the answer choices are ... A.) C(x)=550x -1200 B.) C(x)= 550x + 1200 C.) C(x)=1200x-550 D.) C(x)= 1200x+550

OpenStudy (anonymous):

the answer choices are ... A.) C(x)=550x -1200 B.) C(x)= 550x + 1200 C.) C(x)=1200x-550 D.) C(x)= 1200x+550